The CL.STY Style File

Stuart M. Shieber^{∗}  Julia Hirschberg^{†}
Harvard University  AT&T Bell Laboratories

This sample paper explains and exemplifies how to use the Computational Linguistics L^{A}TEX article style, called “cl.sty”. Most cl.sty commands are the same as with other L^{A}TEX style files (e.g., article.sty). However, there are several new environments and some special instructions are given.

Introduction

This sample paper explains and exemplifies how to use the Computational Linguistics L^{A}TEX article style, called “cl.sty”. Most cl.sty commands are the same as with other L^{A}TEX style files (e.g., article.sty). However, there are several new environments and some special instructions are given.

1. Basics

1.1 Invoking the CL style The file cl.sty is intended to be invoked as a documentstyle option for the document style article. In addition, the twoside document style option should also be used, to generate separate even- and odd-side pages. Your L^{A}TEX file should thus begin with

    \documentstyle[twoside,cl]{article}

Papers published in Computational Linguistics are set in Palatino font. If you have this Postscript font, you can use it by adding the palatino and psgreek style options:

    \documentstyle[twoside,palatino,psgreek,cl]{article}

For information about obtaining the requisite style files, see Section 4.

1.2 Paper-specific information Certain information specific to the paper, like the title and author, is specified with special commands in the preamble. The \title command is used to specify the title, e.g.,

    \title{The CL.STY Style File}

Author information is specified with the \author macro, whose argument is like the argument to a tabular environment. A second line is added for the affiliation, and the \affil macro is used to typeset this information appropriately. Finally, the thanks macro is used to provide additional address information for each author.

    \author{ Stuart M. Shieber%
             \thanks{Division of Applied Sciences,
                     Cambridge, MA 02138}
           & Julia Hirschberg%
             \thanks{600 Mountain Avenue,
                     Murray Hill, NJ 07974} \\

∗ Division of Applied Sciences, Cambridge, MA 02138 † 600 Mountain Avenue, Murray Hill, NJ 07974 \begin{matrix}c \\ ◯ \end{matrix} 1994 Association for Computational Linguistics

Computational Linguistics Volume 16, Number 1

             \affil{Harvard University}
           & \affil{AT\&T Bell Laboratories}}

Information about the issue in which the paper is to appear is declared with the \issue command, which specifies the volume, number, and year of the issue.

    \issue{16}{1}{1994}

Finally, information to be used in the running heads is provided with the \running-

title and \runningauthor commands.
    \runningtitle{The CL.STY Style File}
    \runningauthor{Shieber and Hirschberg}

A footer on the first page can contain copyright information or other publication information. For publication in Computational Linguistics, the footer defaults to an appropriate copyright notice, as seen on the first page. This footer can be overridden by redefining the macro \footmark in the preamble, e.g.,

    \renewcommand{\footmark}{\copyright\ 1894 Association for
                             Computational Phrenology}

(The \footmark and \issue commands are primarily intended for use by the publishers, though you will want to specify a \issue command with some dummy information as a placeholder.)

After all of these preliminaries are set up, you can start the body of the document as follows:

    \begin{document}
    \maketitle

1.3 Abstracts The “abstract” environment is used for the abstract. For the abstract above, for instance, we specified the following:

    \begin{abstract}
    This sample paper ...
    \end{abstract}

1.4 Sections and subsections The sectioning commands are as follows:

    \section{Primary Section}
    \subsection{Secondary section}
    \subsubsection{Tertiary section}
    \paragraph{Quaternary section}

Section 2 shows the generated output for these commands. Unnumbered sections (typically for an introduction, as above), are generated with the \section* command.

Note the capitalization conventions used in these headings. Only primary sections have each word capitalized.

1.5 End matter After the body of the paper, the acknowledgments and references are placed. First, a \starttwocolumn command is used to start two-column layout. Then the acknowledgments environment can be used to provide acknowledgments if any.

The references, inside the thebibliography environment, are set in paragraph form with a small hanging indent. To even off the two columns on the last page, you might add a \newpage to force material to the next column.

The citation commands available in cl.sty are designed to work with the fullname BibTEX bibliography style, which is highly recommended for use with cl.sty. For

Shieber and Hirschberg The CL.STY Style File

Figure 1 The Harvard University shield.

This shield is included because it is completely irrelevant to the topic at hand. Note that the multi-sentence explanatory portion of the caption is given as a separate paragraph from the initial noun-phrase portion. This is achieved by inserting a \captionpar command between the parts.

Table 1 Some important tabular information.

1.1001  2.2220  3.0303  1.1001  2.2220  3.0303
Simple  Complex  Difficult  Simple  Complex  Difficult
NATA  NATB  NATC  NATA  NATB  NATC
dd.f  —  dd.c  dd.f  —  dd.c

further information on this style and the citation macros, see the documentation in

fullname-doc.tex.

2. Primary Section

This section and its various subsections are provided merely to provide an example of the section headings at different levels of embedding.

2.1 Secondary section This section and its various subsections are provided merely to provide an example of the section headings at different levels of embedding.

2.1.1 Tertiary section. This section and its various subsections are provided merely to provide an example of the section headings at different levels of embedding.

Quaternary section. This paragraph is provided merely to provide an example of the section headings at different levels of embedding.

3. Special Environments

The cl.sty style file defines several new environments. This section describes them briefly.

3.1 Figures and tables Figures and tables are placed with the standard L^{A}TEX figure and table environments, as exemplified by Figure 1 and Table 1. Captions, however, appear a bit nonstandard, as seen in Figure 1 above. Figure captions use the \caption command whereas tables use the \tcaption command.

3.1.1 Incorporating Postscript figures. To incorporate Postscript figures, you should use the psfig macros, as was done in this example file. The Postscript file veritas.eps was loaded in Figure 1 as follows:

    \begin{figure}
    \mbox{\psfig{figure=veritas.eps,height=1in}}
    \caption{The Harvard University shield...}
    \label{fig:example}
    \end{figure}

The psfig macros have provision for scaling and rotating figures in various ways. Please see the psfig documentation for further information.

3.2 Itemized lists Itemized lists are specified in the normal way, with the itemize environment. They appear with bullet symbols rather than numbers. Here is an example.

    \begin{itemize}
      \item The first item of the list.
      \item Long items get wrapped and indented appropriately.  The
            appearance of such a long item is exemplified here.  Note
            the attractive layout.
      \begin{itemize}
        \item Items can have subitems.
        \item Perhaps even more than one.
      \end{itemize}
    \end{itemize}

• The first item of the list.

• Long items get wrapped and indented appropriately. The appearance of such a long item is exemplified here. Note the attractive layout.

— Items can have subitems.

— Perhaps even more than one.

Numbered lists are specified with the enumerate environment. Substituting “enumerate” for “itemize” in the above example leads to the following list:

1.The first item of the list.

2.Long items get wrapped and indented appropriately. The appearance of such a long item is exemplified here. Note the attractive layout.

(a)Items can have subitems. (b)Perhaps even more than one.

Similarly, lists with no numbers or bullets can be generated with the outline environment.

The first item of the list.

Long items get wrapped and indented appropriately. The appearance of such a long item is exemplified here. Note the attractive layout.

Items can have subitems. Perhaps even more than one.

3.3 Theorems and theorem-like environments Several theorem-like environments are allowed. Some are given numbers when typeset, e.g., examples, rules, and theorems. Others are unnumbered, e.g., definitions, proofs, and unnumbered theorems. Here are some examples, each followed by a template showing how the environment is specified.

Definition A context-free grammar is a quadruple G = ⟨V, Σ, P, S⟩ where

• V is the vertical rise of the staircase,

• Σ is the number of bartenders that it takes to screw in a light bulb,

• P is the ratio of V to Σ, and

• S is your eldest brother’s hair color.

    \begin{definition}
       ...
    \end{definition}

Example 1 The following grammar is the smallest one specifiable without using the letter ‘b’: ⟨12′, 5, 2.4, brown⟩

    \begin{example}
       ...
    \end{example}

Rule 1 Neither a borrower nor a lender be, unless your name is Morgan, in which case, lender is highly recommended.

    \begin{arule}
       ...
    \end{arule}

Theorem The ratio of the circumference to the diameter is constant.

    \begin{theorem}
       ...
    \end{theorem}

Proof If it weren’t, I’d be likely to know about it, now wouldn’t I. That proves it. Hah. I run rings around you logically.

    \begin{proof}
       ...
    \end{proof}

Theorem 1 The ratio of the circumference to the diameter is even bigger than it was last time I specified a theorem.

    \begin{numtheorem}
       ...
    \end{numtheorem}

3.4 Some more environments Extracts and quotations should use the ext environment, rather than the quote environment. This is specified by

    \begin{ext}
       ...
    \end{ext}

I come now to a preliminary consideration of the question: ‘what is a word?’ But what I have to say now will be supplemented by a detailed discussion at later stages. Words, from the earliest times of which we have historical records, have been objects of superstitious awe. (Russell, 1940, page 21)

For reference — but don’t try this at home — had this been typeset with the quote environment, it would have appeared instead as

I come now to a preliminary consideration of the question: ‘what is a word?’ But what I have to say now will be supplemented by a detailed discussion at later stages.

Words, from the earliest times of which we have historical records, have been objects of superstitious awe. (Russell, 1940, page 21)

Several other environments are defined that may be useful in special circumstances. These include:

• wf: The word formula environment.

• dialog: The dialog environment.

• code: The code environment. (Use of the verbatim environment is typically better for code, however.)

These environments are primarily for use by the publisher; you will probably not need them. However, you can experiment with them to determine if they provide a useful format for your purposes.

4. Obtaining Style Files

The sources for cl.sty and for this documentation are available electronically, as are several style files discussed above for use in preparing papers for Computational Linguistics:

• cl.sty: The L^{A}TEX style for Computational Linguistics articles.

• cl-doc.tex: Documentation on the cl.sty L^{A}TEX style, which you are reading now.

• palatino.sty, psgreek.sty: Style files that cause Postscript Palatino fonts to be used, as is standard for Computational Linguistics articles.

• psfig.tex: A set of macros that allow Postscript figures to be imported into L^{A}TEX documents.

• fullname.bst: A BibTEX style file that is the standard for Computational Linguistics.

• fullname-doc.tex, fullname-doc.bib: Documentation on the fullname

BibTEX style.

These are available by the following methods

• Anonymous ftp from the Association for Computational Linguistics at

“ftp://ftp.cs.columbia.edu/acl-l/Styfiles/” and from the Computation

and Language E-Print Archive (cmp-lg) at

“ftp://xxx.lanl.gov/cmp-lg/papers/macros/''.

• HTTP access from the the Computation and Language E-Print Archive

(cmp-lg) at “http://xxx.lanl.gov/ftp/cmp-lg/papers/macros/”.

Please do not make changes to the style file. If you would like to make changes to some section of the macros (for example, the macros that define the theorems), copy them and put them into a new TEX file named after yourself, e.g., smith.sty. Then read in your new style file when TEXing your article:

\documentstyle[twoside,palatino,psgreek,cl,smith]{article}

Make sure you read your style file in after cl.sty.

5. Conclusion

The cl.sty style file is recommended for preparation of papers for publication in Computational Linguistics.

Acknowledgments The authors wish to thank their mothers. The second author would like further to thank her cat Leo.

References

Andrews, Peter B. 1986. An Introduction to Mathematical Logic and Type Theory: To Truth Through Proof. Computer Science and Applied Mathematics, volume 3, number 2. Orlando, Florida: Academic Press.

Bratman, M. 1983. Castaneda’s theory of thought and action. In James Tomberlin, editor, Agent, Language, and the Structure of the World: Essays Presented to Hector-Neri Castaneda with his Replies. Hackett, Indianapolis, Indiana, pages 149–169.

Dahl, Veronica. 1981. Translating Spanish into logic through logic. Computational Linguistics, 7(3):149–164.

Gazdar, G., E. Klein, G. K. Pullum, and I. A. Sag. 1985. Generalized Phrase Structure Grammar. Oxford: Blackwell Publishing. Also published by Harvard University Press, Cambridge, MA. Gazdar, Gerald, Geoffrey K. Pullum, and Ivan A. Sag. 1982. Auxiliaries and related phenomena in a restrictive theory of grammar. Language, 58:591–638. Russell, Bertrand. 1940. An Inquiry into Meaning and Truth. Allen and Unwin.